OpenStudy (anonymous):

The function models the height y in feet of a stone t seconds after it is dropped from the edge of a vertical cliff. y=-16t^2 + 486 How long will it take for the stone to hit the ground?

OpenStudy (jdoe0001):

have you covered parabolas yet?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

well... notice on this one the leading term is negative meaning the parabola opens downward  so you're really asked what is "t" or the time, when the parabola touches the x-axis or \(\bf y=-16t^2 + 486\implies 0=-16t^2 + 486\qquad \textit{what is }t?\)

OpenStudy (anonymous):

ohhhh ok I get that part I'm confused about what to do next @jdoe0001

OpenStudy (anonymous):

This quadratic equation is similar to a linear one except that now you will have to two values for t.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

BTW we are not here to provide you answers, we are here to help YOU get the answer, and clarify things.

OpenStudy (anonymous):

yes i know. i got an answer, but it is not even close to any of the four answer choices provided

OpenStudy (anonymous):

can you tell us what you did, so we can find the mistake and correct you.

OpenStudy (anonymous):

I used -16t^2 + 486 = 0 then factored that to: -2(8t^2 - 243) = 0

OpenStudy (anonymous):

okay dont factor out 2. just move the 486 to the other side, and treat it like a linear equation. except that you will have a negative and positive answer, since you are going to be square rooting it.

OpenStudy (anonymous):

so just -16t^2 = -486

OpenStudy (anonymous):

yup, isolate for t^2 now. and then square root the both side to get just t=the answer.

OpenStudy (anonymous):

so id divide by -16 on both sides?

OpenStudy (anonymous):

yup

OpenStudy (jdoe0001):

\(\bf 0=-16t^2+486\implies -486=-16t^2\implies \cfrac{\cancel{ -486 }}{\cancel{ -16 }}=t^2 \\ \quad \\ \sqrt{\cfrac{\square }{\square }}=t\)

OpenStudy (jdoe0001):

well.. \(\bf 0=-16t^2+486\implies -486=-16t^2\implies \cfrac{\cancel{ -486 }}{\cancel{ -16 }}=t^2 \\ \quad \\ \pm \sqrt{\cfrac{\square }{\square }}=t\)

OpenStudy (anonymous):

so t^2 = 30.375

OpenStudy (anonymous):

t=5.51 got it

OpenStudy (jdoe0001):

what answer choices did you get?

OpenStudy (anonymous):

Oh and never choose the negative answer for time, since you cannot have a negative time.

OpenStudy (anonymous):

5.51 was one of them

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

k